OpenStudy (anonymous):

If cos theta = -2/3, which of the following are possible? (More than one can be chosen) (will give medal & fan) (listing answers in comments)

OpenStudy (anonymous):

A. sin theta = sqrt(5)/3 and tan theta = sqrt(5)/2 B. sin theta = - sqrt(5)/3 and tan theta = 2/sqrt(5) C. csc theta = -3/2 and tan theta = sqrt(5)/2 D. csc theta = 3/sqrt(5) and tan theta = - sqrt(5)/2

OpenStudy (anonymous):

since \(\sin^2(x)+\cos^2(x)=1\) solve \[b^2+\left(\frac{2}{3}\right)^2=1\] for \(b\)

OpenStudy (anonymous):

or you could go right to \[\sin(x)=\pm\sqrt{1-\left(\frac{2}{3}\right)^2}\]

OpenStudy (anonymous):

I don't know what any of this means. My teacher failed to teach me any of this, so....

OpenStudy (anonymous):

I usually don't ask for a straight up answer, but I'm seriously fed up with all of this.

OpenStudy (anonymous):

ok i will provide an answer

OpenStudy (anonymous):

sine and cosine are points on the unit circle, whose equation is \(a^2+b^2=1\) you have one of them at \(-\frac{2}{3}\) so we can solve \[\left(-\frac{2}{3}\right)^2+b^2=1\] by squaring to get \[\frac{4}{9}+b^2=1\\ b^2=1-\frac{4}{9}\\ b^2=\frac{5}{9}\\ b=\pm\sqrt{\frac{5}{9}}\\ b=\pm\frac{\sqrt5}{3}\]

OpenStudy (anonymous):

all the steps are there, it is not terribly complicated, hope it is more clear

OpenStudy (anonymous):

thank you